Question 13 (1 point) Saved Suppose that all firms in a perfectly competitive market are identical and have the following cosFirms in the short run market equilibrium from question 14 make positive profit. so, eventually new firms will enter the market and sunk fixed cost become avoidable fixed cost and the market enter a new long run market equilibrium.How many firms will exist in this new long run market equilibrium? no units , no roundingQuestion 14 (1 point) Saved Consider the long-run market equilibrium in Question 13 as a starting point. Now suppose that dem
Answer #1
  • Aggregate surplus is the sum of the areas of consumer surplus and poducers surplus.
  • perfect competition maximises total surplus.producing less or more than the comepetitive output lowers total surplus.
  • At the competitive equilibrium e1, with Q1 and p1

TS=A+B+C+D+E.

  • producing less at e2 Q2 and p2

TS2=A+B=D.TS2 less than TS1

  • As a consequences of roducing less C+E are lost.
  • C+E is the deadweight loss [DWL]
  • DWL is the net reduction in total surplus from a loss of surplus by one group that is not offset by a gain to another group from an action that alters a market equilibrium. Explained with the help of diagram given below:p, $ per unit Supply р \ Р Epal Demand MC Q2 Q1 Q, Units per year

Competive output Smaller Output Change {2-1}

Consumer Surplus (CS}

producer Surplus {ps}

A+B+C

D+E

A

B+D

-B-C=^CS

B-E=^ps

Total Surplus {TS}=cs+ps A|+B+C+D+E A+B+D -C-E=^TS=DWL